OpenStudy (anonymous):

e^log3 - e^log 2 Please explain aswell thanks :)

OpenStudy (anonymous):

Let y=e^log3-e^log 2 ln y=ln e^log3-ln^log2 ln y=log3-log2 ln y=log(3/2)

OpenStudy (anonymous):

the answer at the back of the book is one lol

OpenStudy (anonymous):

What is the question, or full instructions. I just took a stab at it.